Mrs. Valle is five times as old as her son. Three years ago, her age was eight times that of her son's age then. Find each of their ages right now.
Answer by stanbon(75887)

Mrs. Valle is five times as old as her son.
Three years ago, her age was eight times that of her son's age then.
Find each of their ages right now.
------------------
Equations:
v = 5s
(v-3) = 8(s-3)
----
Substitute and solve for "s":
5s-3 = 8s-24
3s = 21
s = 7 (son's age now)
----
Solve for "v":
v = 5s
v = 5*7 = 35 (Mrs. Valle's age now)
====================================
Cheers,
Stan H.
====================================